OpenStudy (anonymous):

the great circle of has an area of 400 pie ft ^2. Find the surface area and the volume

OpenStudy (anonymous):

we have to first find radius r of the sphere using the of the great circle since area =pi(r^2) thus r^2 =(400 pi)/(pi) hece r^2=400 =20x20 hence r=20 ft now u can use the formula for volume =(4/3)(pi)(r^3) (use r=20 here ) again for surface area =4(pi)(r^2) (here also use r=20) i hope mere calculation will help u to get the result